# The Experimental Unit Information Index: Balancing Evidentiary Value and Sample Size of Adaptive Designs

### v1

@@ Abstract:
Reducing the number of experimental units is one of the three pillars of the 3R principles (Replace, Reduce, Refine) in animal research. At the same time, statistical error rates need to be controlled to enable reliable inferences and decisions. This paper proposes a novel measure to quantify the evidentiary value of one experimental unit for a given study design. The experimental unit information index (EUII) is based on power, Type-I error and sample size, and has attractive interpretations both in terms of frequentist error rates and Bayesian posterior odds. We introduce the EUII in simple statistical test settings and show that its asymptotic value depends only on the assumed relative effect size under the alternative. We then extend the definition to adaptive designs where early stopping for efficacy or futility may cause reductions in sample size. Applications to group-sequential designs and a recently proposed adaptive statistical test procedure show the usefulness of the approach when the goal is to maximize the evidentiary value of one experimental unit.
@@ 1 Introduction
Over the last decades, various adaptive designs have been developed where an experiment can be stopped early, either for efficacy (a convincing result has already been obtained) or for futility (the results are so unconvincing that there is only little sense in continuing the experiment). Stopping an experiment early ensures that the number of animals included in an experiment does not become too large. This reflects the second pillar of the 3R principles “Replace, Reduce, Refine” in animal research (Russell and Burch, 1959), described as
@@
to use methods for obtaining comparable levels of information from the use of fewer animals in scientific procedures, or for obtaining more information from the same number of animals” (swiss3rcc.org/3rs-for-the-public/)
@@
The 3R principles are now widely adopted to improve the design and analysis of animal studies (Reynolds, 2023, 2024a, 2024b; Townsend et al., 2025). However, the focus of standard statistical theory on error rates ignores that different adaptive designs may have different (expected) sample sizes. Furthermore, a comparison of methods with desired Type-I error rate control with alternative methods that may have more power but also Type-I error rate inflation can be difficult. To address the latter issue, optimising the trade-off between Type-I and Type-II error rates has been recently considered Mudge et al. (2012); Walley and Grieve (2021); Grieve (2024). The goal is here to optimize a weighted sum of error rates based on prior probabilities of the null and the alternative hypothesis and the relative costs of Type-I and Type-II errors. An alternative approach is to adjust power for the corresponding Type-I error rate using the theory of receiver operating characteristic curves (Lloyd, 2005; Cavus et al., 2019). However, all these methods do not take sample size into account.
@@
In this paper we propose a new measure of statistical information that takes into account Type-I error rate, power, and sample size. This measure is related to the diagnostic odds ratio to quantify diagnostic accuracy (Glas et al., 2003), which is widely used in clinical epidemiology and which also has attractive interpretations when applied to statistical hypothesis tests. First, it has a frequentist interpretation as the ratio of the rejection odds under the alternative and under the null hypothesis. Second, it also has a Bayesian interpretation as the ratio of the posterior odds of the alternative hypothesis given a significant versus nonsignificant results. Finally, the diagnostic odds ratio allows for a normalization with respect to the (expected) sample size to obtain a measure of the evidentiary value of one unit/animal/patient, the experimental unit information index (EUII).
@@
This article is organized as follows: Section 2 reviews likelihood ratios and diagnostic odds ratios and applies them to statistical hypothesis tests. Section 3 then defines the experimental unit information index for fixed sample size and studies its performance in standard statistical test settings. The extension to adaptive tests with different expected sample size of significant and nonsignificant findings is described in Section 4. Applications to group-sequential designs (Section 4.1) and “constrained sample augmentation” (Section 4.2), an adaptive test procedure recently proposed by Reinagel (2023), are described. We close with some discussion in Section 5.
@@ 2.1 Likelihood ratios and the diagnostic odds ratio
In a diagnostic context, $H_{1}$ represents the hypothesis that a patient has the disease whereas $H_{0}$ represents that the patient does not have the disease. Likelihood ratios (LRs) are commonly used in diagnostic testing to quantify the (prior to posterior) change in odds $\mathsf{Odds}(H_{1})=\Pr(H_{1})/\Pr(H_{0})$ after having observed a positive or negative test result (Deeks and Altman, 2004):
@@
Useful diagnostic tests have positive likelihood ratio $\mbox{LR}^{+}>1$ and negative likelihood ratio $\mbox{LR}^{-}<1$, where one of the two inequalities implies the other one. Equation (5) shows that a positive result from a useful test increases the odds (and therefore also the probability) of having the disease whereas a negative result decreases it. The diagnostic odds ratio is thus a single indicator of diagnostic test accuracy, the larger, the better. We will describe its properties in more detail in the following, when we apply it to statistical hypotheses tests.
@@ 3 The experimental unit information index
The motivation to take the $n$th root comes from equation (8) and (9), respectively. These two equations show that likelihood ratios act multiplicatively on the odds of $H_{1}$, transforming prior odds to posterior odds. Both likelihood ratios are based on data from $n$ experimental units, so evey unit contributes on average an amount of $({\mbox{LR}^{-}})^{1/n}$ and $({\mbox{LR}^{+}})^{1/n}$, respectively, to the likelihood ratios, because
@@ 3.1.1 Standard z-test
Table 1 gives the DOR and the EUII also for other values of the power and Type-I error rate. We can see that a test with 90% power at significance level 5% has more evidentiary value than a test with 80% power at significance level 2.5% (DOR = 171 vs. 156). The increase in DOR is parallelized by an increase in the required sample size $n$ to detect the standardized effect $\delta=0.5$, so the corresponding values of EUII are nearly the same, 1.13 and 1.14, respectively. The EUII takes its largest value 1.15 if both Type-I and Type-II error have the largest value (5% and 20%, respectively). We will see later, that the asymptotic value of EUII (for fixed T1E rate and increasing sample size $n$) is $\exp(\delta^{2}/2)=1.133$, so the finite sample size values of EUII in Table 1 are all close to this asymptotic value.
Table 1: Diagnostic odds ratio (DOR) and experimental unit information index (EUII) for different values of power and Type-I error rate. Also shown is the required sample size $n$ to detect a standardized effect of $\delta=0.5$ with a one-sample $z$-test and the corresponding value of EUII.
@@
If the design would have been instead based on two equally-sized samples and $\delta$ represents the standardized difference to be detected, then the required total sample size $n=125.6$ at the 5% significance level and 80% power is four times as large (Kirkwood and Sterne, 2003), and we obtain
@@
The average evidentiary increase per unit is now 3.5%. Note that $\mbox{EUII}_{1}=({\mbox{EUII}_{2}})^{4}$ which means that one unit in a one-sample study contributes as much information as four units in a (perfectly balanced) two-sample trial, if the goal is to detect the same standardized effect.
@@ 3.2 Z-test and t-test with increasing sample size
Consider now the standard one-sample one-sided $t$-test to test the null hypothesis $H_{0}\colon\mu=0$ versus the alternative $H_{1}\colon\mu>0$. We let the sample size $n$ vary between $2^{3}=8$ and $2^{13}=8192$ and compute the power to detect a standardized effect size $\delta$ at significance level $\alpha=5\%$. Figure 2 shows how the resulting EUII varies with $\delta$ and $n$. Not surprisingly, EUII increases with increasing $\delta$, because the power to detect $\delta$ increases with increasing $\delta$ (for fixed $n$). If $\delta=0.1$ and $n=8$, EUII is smaller than 1 because the power to detect such a small effect with such a small sample size is in fact smaller than the T1E rate. Such a test is called biased in the mathematical statistics literature (Lehmann, 1959).
Figure 2: The experimental unit information index for the standard one-sample one-sided $t$-and $z$-test as a function of sample size and effect size. The asymptotic value $\exp(\delta^{2}/2)$ for the different values of the standardized effect size $\delta$ is indicated by the dotted lines.
@@ 4 The experimental unit information index for adaptive designs
In adaptive trial designs, a trial can often be stopped preliminary if the results are sufficiently compelling at an interim analysis, for example, if there is strong evidence for treatment benefit or harm (Kairalla et al., 2012). This means that the sample size is random and typically differs under $H_{0}$ and $H_{1}$. In the following, we denote a random sample size by $N$ to distinguish it from a fixed sample size $n$. An adaptive trial can then be characterized by the pair of the sample size $N$ at which the trial stops and the indicator $S$ of whether the results were (non)significant. The expected sample sizes for significant (++) and nonsignificant (-) are then then defined by the conditional expectations $\operatorname{\mathsf{E}}(N_{+})=\operatorname{\mathsf{E}}(N\,|\,S=\text{significant})$ and $\operatorname{\mathsf{E}}(N_{-})=\operatorname{\mathsf{E}}(N\,|\,S=\text{nonsignificant})$, respectively.
@@
The rationale for the generalization (18) is that the computation of posterior predictive odds with likelihood ratios is either based on significant (equation (8)) or nonsignificant findings (equation (9)). It is therefore natural to normalize the numerator $\mbox{LR}^{+}$ in (18) with the expected sample size of significant findings $\operatorname{\mathsf{E}}(N_{+})$ and the denominator with the expected sample size of nonsignificant findings $\operatorname{\mathsf{E}}(N_{-})$.
@@ 4.1 Group-sequential designs
Note that Pocock and O’Brien-Fleming exactly control the Type-I error rate at $\alpha=2.5\%$ whereas the one-sided Haybittle-Peto method (Blenkinsop et al., 2019) (with nominal thresholds $a_{1}=a_{2}=a_{3}=0.0005$ and $a_{4}=0.025$) has a slightly increased overall T1E rate of 2.54%. The maximum sample size $n_{\scriptsize\texttt{Max}}$ (last column of Table 3) from Pocock and O’Brien-Fleming is larger than the corresponding sample size 50 under a fixed design with the same Type-I error rate and power of 90% (corresponding to a standardized effect size of $\delta=0.46$). Without futility stopping, this is also the sample size for any nonsignificant result, both under the null and the alternative: $\operatorname{\mathsf{E}}(N_{0-})=\operatorname{\mathsf{E}}(N_{1-})=\operatorname{\mathsf{E}}(N_{-})=n_{\scriptsize\texttt{Max}}$. However, the expected sample size for significant results will be smaller and in general depends on the underlying hypothesis and the assumed power. The maximum sample size $n_{\scriptsize\texttt{Max}}=50$ from Haybittle-Peto is chosen to be equal to the sample size from the fixed design. Although the power and Type-I error rate of Haybittle-Peto does not correspond in this case to the other methods, the EUII enables a meaningful comparison.
Figure 3 displays the EUII for the different methods (top), relative to the fixed design (middle), and the maximum sample size and expected sample size of significant findings (bottom) for values of the standardized effect size $\delta$ between 0.1 and 0.8. As expected, the EUII increases with increasing $\delta$ but also depends on the assumed prior probability of $H_{1}$, chosen to be either 0.01, 0.1 or 0.5. In particular, for $\Pr(H_{1})=.01$ the EUII depends less on the effect size $\delta$ under the alternative, because the second term $\operatorname{\mathsf{E}}(N_{1+})$ (that depends on $\delta$) of the expected sample size (20) for significant findings receives less weight.
@@
Figure 3 also shows the expected sample size of significant findings under the two hypotheses, for the different values of the standardized effect size. With the exception of Pocock, the expected sample size of significant findings is always larger under $H_{0}$ than under $H_{1}$. Only for Pocock it can be the other way round if the standardized effect size is less than $\delta=0.46$, which corresponds to the power being below 90%. This is a consequence of the fact that, under $H_{0}$, a significant finding with Pocock’s method is more likely to be obtained at the first interim analysis and less at the second and so forth. This is not the case under $H_{1}$, where a significant finding at the second interim analysis is more likely than at the first. This peculiar feature is caused by the constant and relatively large nominal significance levels (see Table 3) and has been one of the criticisms of the Pocock method in clinical trials, where – under the $H_{0}$ – it is less difficult that a trial is stopped early than for competing methods (Simon, 1994; Matthews, 2006). However, this feature has been described as an adventage in animal trials where the goal is to minimize animal numbers (Blotwijk et al., 2022) and our analysis confirms that smaller expected sample sizes under $H_{0}$ lead to a larger evidentiary value per animal of the Pocock method compared to the other ones. For example, the EUII is smaller for Haybittle-Peto, where the nominal significance levels (with the exception of the last one) are also constant, but smaller. Significant findings are then more likely to be obtained at the last analysis and the expected sample size under $H_{0}$ increases accordingly. In sum, the EUII allowed us to compare sequential designs with differing power, Type-I error rate, sample sizes, and interim analyses time points.

### v2

# Balancing Evidentiary Value and Sample Size of Adaptive Designs with Application to Animal Experiments

@@ Abstract:
Reducing the number of experimental units is one of the three pillars of the 3R principles (Replace, Reduce, Refine) in animal research. At the same time, statistical error rates need to be controlled to enable reliable inferences and decisions. This paper proposes a novel measure to quantify the evidentiary value of one experimental unit for a given study design. The experimental unit information index (EUII) is based on power, Type-I error and sample size, and has attractive interpretations both in terms of frequentist error rates and Bayesian posterior odds. We introduce the EUII in simple statistical test settings and show that its asymptotic value depends only on the assumed relative effect size under the alternative. We then extend the definition to adaptive designs where early stopping for efficacy or futility may cause reductions in sample size. Applications to group-sequential designs and a recently proposed adaptive statistical test procedure show the usefulness of the approach when the goal is to maximize the evidentiary value of one experimental unit. A reanalysis of 2738 animal experiments with simulated results from (post-hoc) interim analyses illustrates the possible savings in sample size.
@@ 1 Introduction
Over the last decades, various adaptive designs have been developed where an experiment can be stopped early, either for efficacy (a convincing result has already been obtained) or for futility (the results are so unconvincing that there is only little sense in continuing the experiment). Stopping an experiment early ensures that the number of animals included in an experiment does not become too large. These questions are of clear practical relevance in animal research, as exemplified by the recently proposed Stroke Preclinical Assessment Network (SPAN) experimental design (Lyden et al., 2022), calling for more efficient and rigorous designs in preclinical research with stops for both futility and efficacy. This also reflects the second pillar of the 3R principles “Replace, Reduce, Refine” in animal research (Russell and Burch, 1959), described as
@@
to use methods for obtaining comparable levels of information from the use of fewer animals in scientific procedures, or for obtaining more information from the same number of animals” (swiss3rcc.org/3rs-for-the-public/).
@@
The 3R principles are now widely adopted to improve the design and analysis of animal studies (Reynolds, 2023, 2024a, 2024b; Townsend et al., 2025). However, the focus of standard statistical theory on statistical significance ignores that different adaptive designs may have different (expected) sample sizes. Furthermore, a comparison of methods with desired Type-I error rate control with alternative methods that may have more power but also Type-I error rate inflation can be difficult. To address the latter issue, optimising the trade-off between Type-I and Type-II error rates has been recently proposed (Mudge et al., 2012; Walley and Grieve, 2021; Grieve, 2024). The goal is here to optimise a weighted sum of error rates based on prior probabilities of the null and the alternative hypothesis and the relative costs of Type-I and Type-II errors. An alternative approach is to adjust power for the corresponding Type-I error rate using the theory of receiver operating characteristic curves (Lloyd, 2005; Cavus et al., 2019). However, all these methods do not take sample size into account.
To put the discussion into context, consider results from 2738 animal experiments in neuroscience and metabolism, collected by Bonapersona et al. (2021) through a systematic search of studies that were previously included in meta-analyses published between 2003 and 2019. Figure 1 shows the distribution of sample sizes and standardised mean difference effect sizes (Hedges’ $g$). Most studies have total sample sizes between 10 and 30 animals but there are some studies with more than 100 animals. The majority of standardised mean differences are below 1 in absolute size, but some studies have remarkably large effect sizes. This raises the question how many animals could have been saved if the experiments had incorporated a sequential design allowing for interim analyses. A related question is which type of group-sequential design (Jennison and Turnbull, 1999) should be used to minimise the expected costs and sample sizes of the experiments while ensuring that the chosen design has sufficiently high evidentiary value as quantified by statistical error rates (Blotwijk et al., 2022).
Figure 1: Distribution of sample sizes and standardised mean differences in 2738 animal experiments in neuroscience and metabolism (Bonapersona et al., 2021). Extreme standardised mean differences with absolute value larger than 5 are not shown ($n=28$).
@@
In this paper we propose a new measure of statistical information that takes into account Type-I error rate, power, and sample size. This measure is related to the diagnostic odds ratio to quantify diagnostic accuracy (Glas et al., 2003), which is widely used in clinical epidemiology and has attractive interpretations when applied to statistical hypothesis tests. First, it has a frequentist interpretation as the ratio of the rejection odds under the alternative and under the null hypothesis. Second, it also has a Bayesian interpretation as the ratio of the posterior odds of the alternative hypothesis given a significant versus nonsignificant results. Finally, the diagnostic odds ratio allows for a normalisation with respect to the (expected) sample size to obtain a measure of the evidentiary value of one unit/animal/patient, the experimental unit information index (EUII).
@@
This article is organized as follows: Section 2 reviews likelihood ratios and diagnostic odds ratios and applies them to statistical hypothesis tests. Section 3 then defines the experimental unit information index (EUII) for fixed sample size and studies its performance in standard statistical test settings. The extension to adaptive tests with different expected sample size of significant and nonsignificant findings is described in Section 4. Applications to group-sequential designs (Section 4.1) and “constrained sample augmentation” (Section 4.2), an adaptive test procedure recently proposed by Reinagel (2023), are described. In a simulation study we compare the EUII of constrained sample augmentation with group-sequential methods with additional stops for futility based on predictive power. Based on our findings, we re-analyse the Bonapersona et al. (2021) data set in Section 5 to assess how many animals could have been saved with group-sequential designs, and close with some discussion in Section 6.
@@ 2.1 Likelihood ratios and the diagnostic odds ratio
In a diagnostic context, $H_{1}$ represents the hypothesis that a patient has the disease whereas $H_{0}$ represents that the patient does not have the disease. Likelihood ratios (LRs) are commonly used to quantify the (prior to posterior) change in odds $\mathsf{Odds}(H_{1})=\Pr(H_{1})/\Pr(H_{0})$ after having observed a positive or negative test result (Deeks and Altman, 2004):
@@
Useful diagnostic tests have positive likelihood ratio $\mbox{LR}^{+}>1$ and negative likelihood ratio $\mbox{LR}^{-}<1$, where one of the two inequalities implies the other one. Equation (5) shows that a positive result from a useful test increases the odds (and therefore also the probability) of having the disease whereas a negative result decreases it. The diagnostic odds ratio is thus a single indicator of diagnostic test accuracy, the larger, the better. We will describe its properties in more detail in the following, when we apply it to statistical hypothesis tests.
@@ 3 The experimental unit information index
The motivation to take the $n$th root comes from equation (8) and (9), respectively. These two equations show that likelihood ratios act multiplicatively on the odds of $H_{1}$, transforming prior odds to posterior odds. Both likelihood ratios are based on data from $n$ experimental units, so every unit contributes on average an amount of $({\mbox{LR}^{-}})^{1/n}$ and $({\mbox{LR}^{+}})^{1/n}$, respectively, to the likelihood ratios, because
@@ 3.1.1 Standard z-test
Table 1 gives the DOR and the EUII also for other values of the power and Type-I error rate. We can see that a test with 90% power at significance level 5% has more evidentiary value than a test with 80% power at significance level 2.5% (DOR = 171 vs. 156). The increase in DOR is parallelized by an increase in the required sample size $n$ to detect the standardised effect $\delta=0.5$, so the corresponding values of EUII are nearly the same, 1.13 and 1.14, respectively. The EUII takes its largest value 1.15 if both Type-I and Type-II error have the largest value (5% and 20%, respectively). We will see later, that the asymptotic value of EUII (for fixed T1E rate and increasing sample size $n$) is $\exp(\delta^{2}/2)=1.133$, so the finite sample size values of EUII in Table 1 are all close to this asymptotic value.
Table 1: Diagnostic odds ratio (DOR) and experimental unit information index (EUII) for different values of power and Type-I error rate. Also shown is the required sample size $n$ to detect a standardised effect of $\delta=0.5$ with a one-sample $z$-test and the corresponding value of EUII.
@@
If the design would have been instead based on two equally-sized samples and $\delta$ represents the standardised difference to be detected, then the required total sample size $n=125.6$ at the 5% significance level and 80% power is four times as large (Kirkwood and Sterne, 2003), and we obtain
@@
The average evidentiary increase per unit is now 3.5%. Note that $\mbox{EUII}_{1}=({\mbox{EUII}_{2}})^{4}$ which means that one unit in a one-sample study contributes as much information as four units in a (perfectly balanced) two-sample trial, if the goal is to detect the same standardised effect.
@@ 3.2 Z-test and t-test with increasing sample size
Consider now the standard one-sample one-sided $t$-test to test the null hypothesis $H_{0}\colon\mu=0$ versus the alternative $H_{1}\colon\mu>0$. We let the sample size $n$ vary between $2^{3}=8$ and $2^{13}=8192$ and compute the power to detect a standardised effect size $\delta$ at significance level $\alpha=5\%$. Figure 3 shows how the resulting EUII varies with $\delta$ and $n$. Not surprisingly, EUII increases with increasing $\delta$, because the power to detect $\delta$ increases with increasing $\delta$ (for fixed $n$). If $\delta=0.1$ and $n=8$, EUII is smaller than 1 because the power to detect such a small effect with such a small sample size is in fact smaller than the T1E rate. Such a test is called biased in the mathematical statistics literature (Lehmann, 1959).
Figure 3: The experimental unit information index for the standard one-sample one-sided $t$-and $z$-test as a function of sample size and effect size. The asymptotic value $\exp(\delta^{2}/2)$ for the different values of the standardised effect size $\delta$ is indicated by the dotted lines.
@@ 4 The experimental unit information index for adaptive designs
In adaptive trial designs, a trial can often be stopped preliminary if the results are sufficiently compelling at an interim analysis, for example, if there is strong evidence for treatment benefit or harm (Kairalla et al., 2012). This means that the sample size is random and typically differs under $H_{0}$ and $H_{1}$. In the following, we denote a random sample size by $N$ to distinguish it from a fixed sample size $n$. An adaptive trial can then be characterised by the pair of the sample size $N$ at which the trial stops and the indicator $S$ of whether the results were (non)significant. The expected sample sizes for significant (++) and nonsignificant (-) are then then defined by the conditional expectations $\operatorname{\mathsf{E}}(N_{+})=\operatorname{\mathsf{E}}(N\,|\,S=\text{significant})$ and $\operatorname{\mathsf{E}}(N_{-})=\operatorname{\mathsf{E}}(N\,|\,S=\text{nonsignificant})$, respectively.
@@
The rationale for the generalization (18) is that the computation of posterior predictive odds with likelihood ratios is either based on significant (equation (8)) or nonsignificant findings (equation (9)). It is therefore natural to normalise the numerator $\mbox{LR}^{+}$ in (18) with the expected sample size of significant findings $\operatorname{\mathsf{E}}(N_{+})$ and the denominator with the expected sample size of nonsignificant findings $\operatorname{\mathsf{E}}(N_{-})$.
@@ 4.1 Group-sequential designs
Note that Pocock and O’Brien-Fleming exactly control the Type-I error rate at $\alpha=2.5\%$ whereas the one-sided Haybittle-Peto method (Blenkinsop et al., 2019) (with nominal thresholds $a_{1}=a_{2}=a_{3}=0.0005$ and $a_{4}=0.025$) has a slightly increased overall T1E rate of 2.54%. The maximum sample size $n_{\scriptsize\texttt{Max}}$ (last column of Table 3) from Pocock and O’Brien-Fleming is larger than the corresponding sample size 50 under a fixed design with the same Type-I error rate and power of 90% (corresponding to a standardised effect size of $\delta=0.46$). Without futility stopping, this is also the sample size for any nonsignificant result, both under the null and the alternative: $\operatorname{\mathsf{E}}(N_{0-})=\operatorname{\mathsf{E}}(N_{1-})=\operatorname{\mathsf{E}}(N_{-})=n_{\scriptsize\texttt{Max}}$. However, the expected sample size for significant results will be smaller and in general depends on the underlying hypothesis and the assumed power. The maximum sample size $n_{\scriptsize\texttt{Max}}=50$ from Haybittle-Peto is chosen to be equal to the sample size from the fixed design. Although the power and Type-I error rate of Haybittle-Peto does not correspond in this case to the other methods, the EUII enables a meaningful comparison.
Figure 4 displays the EUII for the different methods (top), relative to the fixed design (middle), and the maximum sample size and expected sample size of significant findings (bottom) for values of the standardised effect size $\delta$ between 0.1 and 0.8. As expected, the EUII increases with increasing $\delta$ but also depends on the assumed prior probability of $H_{1}$, chosen to be either 0.01, 0.1 or 0.5. In particular, for $\Pr(H_{1})=.01$ the EUII depends less on the effect size $\delta$ under the alternative, because the second term $\operatorname{\mathsf{E}}(N_{1+})$ (that depends on $\delta$) of the expected sample size (20) for significant findings receives less weight.
@@
Figure 4 also shows the expected sample size of significant findings under the two hypotheses, for the different values of the standardised effect size. With the exception of Pocock, the expected sample size of significant findings is always larger under $H_{0}$ than under $H_{1}$. Only for Pocock it can be the other way round if the standardised effect size is less than $\delta=0.46$, which corresponds to the power being below 90%. This is a consequence of the fact that, under $H_{0}$, a significant finding with Pocock’s method is more likely to be obtained at the first interim analysis and less at the second and so forth. This is not the case under $H_{1}$, where a significant finding at the second interim analysis is more likely than at the first. This peculiar feature is caused by the constant and relatively large nominal significance levels (see Table 3) and has been one of the criticisms of the Pocock method in clinical trials, where – under the $H_{0}$ – it is less difficult that a trial is stopped early than for competing methods (Simon, 1994; Matthews, 2006). However, this feature has been described as an advantage in animal trials where the goal is to minimize animal numbers (Blotwijk et al., 2022) and our analysis confirms that smaller expected sample sizes under $H_{0}$ lead to a larger evidentiary value per animal of the Pocock method compared to the other ones. For example, the EUII is smaller for Haybittle-Peto, where the nominal significance levels (with the exception of the last one) are also constant, but smaller. Significant findings are then more likely to be obtained at the last analysis and the expected sample size under $H_{0}$ increases accordingly. In sum, the EUII allowed us to compare sequential designs with differing power, Type-I error rate, sample sizes, and interim analyses time points.